Toler Company sells flags with team logos. Toler has fixed costs of $900,000 per year plus variable costs of $10.00 per flag.Requirement 1. Use the equation approach to compute the number of flags Toler must sell each year to break even. First, selecRequirement 1. Use the equation approach to compute the number of flags Toler must sell each year to break even. First, selecRequirement 2. Use the contribution margin ratio approach to compute the dollar sales Toler needs to earn $75,000 in operatinRequirement 2. Use the contribution margin ratio approach to compute the dollar sales Toler needs to earn $75,000 in operatinRequirement 3. Prepare Tolers contribution margin income statement for the year ended December 31, 2018, for sales of 52,000Requirement 3. Prepare Tolers contribution margin income statement for the year ended December 31, 2018, for sales of 52,000Requirement 4. The company is considering an expansion that will increase fixed costs by 30% and variable costs by $2.50 perRequirement 4. The company is considering an expansion that will increase fixed costs by 30% and variable costs by $2.50 per
Answer #1

1.
Sales Revenue - Variable Costs - Fixed Costs = Target Profit
$25x - $10x - $900000 = 0
x = 60000 units to break even

2.
Required Sales in Dollars = (Fixed COsts + Target Profit) / CM Ratio
= ($900000 + 75000) / 60% = $1625000
CM Ratio = 15/25 = 60%

3.

Sales Revenue $   13,00,000
Variable Costs $     5,20,000
Contribution Margin $     7,80,000
Fixed Costs $     9,00,000
Net Operating Income $    -1,20,000

4.
Sales Revenue - Variable Costs - Fixed Costs = Target Profit
$25x - $12.5x - $900000x1.3 = 0
x = 93600 units to break even

Break even point in dollars = 93600 x $25 = $2340000

Toler should only undertake the expansion if expected profits from expansion are greater than expected costs